(1 point) Let u = (-2,-3) and v = (-1,6). Then u+v=< >, u-v=< -3v=< u.V= and || 0 ||
Answer #1

and va 0 = (-2-3) and (9 (1) 316-را- 1- = 7 +0 - 3t 3, 2) ン = لا - ا (3-6- را+ 2 - (9- را) - = لا 3 - (2 - 18) U: V = (۱ ٫6-)